A body of volume 36cc floats with ¾ of its volume submerged in water. The density of body is



a) 0.25 g/cc



b) 0.75 g/cc



c) 0.9 g/cc



d) 0.1 g/cc


1
Expert's answer
2022-02-14T13:56:31-0500

Answer

body of volume V=36cc=36*10-6m

Submerged in water V'=27*10-6 m

Using the buoyancy force

"F_B=W\\\rho_wV'g=\rho Vg\\"

So density of body

"\rho =\frac{\rho_w V'}{V}\\=\frac{1000*27*10^{-6}}{36 *10^{-6} }\\=0.75g/cc"